OpenStudy (anonymous):

i am stuck on this problem and none of the multiple choice answers fit the uk, denmark, and norway have a total of 20 territories. norway has 1/5 as many as the uk. denmark has 2 territories. how many territories does norway have.

OpenStudy (anonymous):

hey, how's it going

OpenStudy (anonymous):

good thanks

OpenStudy (anonymous):

uk denmark and norway have 20 territories, denmark has 2 territories so uk+norway = 18

OpenStudy (anonymous):

now you need to set up a system of equations to solve

OpenStudy (anonymous):

the first one is u+n=18

OpenStudy (anonymous):

i got that then i divided it by .2 but it didn't fit

OpenStudy (anonymous):

try to turn the sentence-form information they give you into a math equation

OpenStudy (anonymous):

divided what by .2? did you understand the idea of a system of equations?

OpenStudy (anonymous):

the 18 because it said that had 1/5 as many territories...sorry meant multiplied 18 by .2

OpenStudy (anonymous):

basically if you have two variables, you need two equations to work with to solve for the variables

OpenStudy (anonymous):

1/5 as many territories as the UK

OpenStudy (anonymous):

not 1/5 as many territories as 18

OpenStudy (anonymous):

opps

OpenStudy (anonymous):

norway has 1/5 as many as the uk.

OpenStudy (anonymous):

norway = 1/5 * uk

OpenStudy (anonymous):

n = 1/5 * u

OpenStudy (anonymous):

sorry we just start algebra and the teacher didn't explain this and i couldnt find an example in the book

OpenStudy (anonymous):

don't be sorry, it's no problem

OpenStudy (anonymous):

now that you have n+u=18 and n=1/5 * u, you want to substitute n back into the first equation so you have only 1 variable to work with

OpenStudy (anonymous):

i'm so confused :(

OpenStudy (anonymous):

would it be 1/5 * u+u=18 or 2u=18-1/5 or something like that

OpenStudy (anonymous):

are you still able to help me

OpenStudy (anonymous):

thanks for your help i will ask the teacher thank you i have to go to bed